OpenStudy (anonymous):

A cylindrical piece of iron pipe is shown below. The wall of the pipe is 1.25 inches thick. What is the approximate inside volume of the pipe? 69 cubic inches 99 cubic inches 332 cubic inches 703 cubic inches

OpenStudy (anonymous):

do we know what the height and radius are?

OpenStudy (anonymous):

oh ya sorry height is 14 inches and diameter is 8

OpenStudy (anonymous):

does the diameter given include the thickness of the cylinder?

OpenStudy (anonymous):

lol I really don't know since that's all the question says XD, but I know its either 332 or 703

OpenStudy (anonymous):

well you know the equation right?

OpenStudy (anonymous):

ya I think we have to find the area of the circle then multiply it by the height

OpenStudy (anonymous):

or the volume of a cylinder formula?

OpenStudy (anonymous):

yes \[\pi \times r ^{2}\times h\]

OpenStudy (anonymous):

thats the equation for the V of a cylinder

OpenStudy (anonymous):

so we don't have to use the thickness? sorry im a bit confused with this chapter XD

OpenStudy (anonymous):

ok so its 3.14*4^2*14?

OpenStudy (anonymous):

well lets just assume that the diameter includes the thickness... we need to subtract the thickness twice to compensate for both sides the equation should be \[\pi \times (\frac{ 10- 2 \times 1.25 }{ 2 })^{2} \times 14\]

OpenStudy (anonymous):

sorry give me a sec to solve lol

OpenStudy (anonymous):

i think your choices are off